OpenStudy (anonymous):

What is the equation of a circle with a center at (0, 0) and a radius of 4.5?

OpenStudy (er.mohd.amir):

(x-h)^2+(y-k)^2=r^2 where (h,k) is center and r is radius

OpenStudy (anonymous):

so (x-0)^2+y-0^2=4.5^2

OpenStudy (anonymous):

@Er.Mohd.AMIR